OpenStudy (anonymous):

if cot of theta = -9/2 with theta in Quadrant 2, find the sin of theta. How do i solve this?

OpenStudy (imstuck):

Oh I love these! I will help you!

OpenStudy (imstuck):

First tell me what the cotangent identity is the same as (like tangent is the same as sine/cos). Give me that first.

OpenStudy (anonymous):

ok i think its cos/sine?

OpenStudy (imstuck):

Yes, that's right. So another way to write that is this\[\cot \theta=-\frac{ 9 }{ 2 }-->\frac{ \cos \theta }{ \sin \theta }=-\frac{ 9 }{ 2 }\]

OpenStudy (imstuck):

In the second quadrant which identity is negative, cos or sin?

OpenStudy (anonymous):

cosine is negative in the second quadrant

OpenStudy (imstuck):

yes. Does cos represent your x value or your y value?

OpenStudy (imstuck):

I wonder if there's a problem with the connection here; it seems as if you have been typing forever. and there's nothing there.

OpenStudy (anonymous):

I believe it is x value

OpenStudy (anonymous):

sorry my computer just froze

OpenStudy (anonymous):

im back now

OpenStudy (imstuck):

Yes, cos is your x value. You're doing great! So let's draw that.

OpenStudy (imstuck):

Your goal here is to find sin of that angle. In order to do that you need to find the hypotenuse, since the sin of an angle = side opposite the angle/hypotenuse, right? So can you find x?

OpenStudy (imstuck):

Just a little light humor.

OpenStudy (imstuck):

What do you need to do to find x?

OpenStudy (anonymous):

yes here let me do it real quick

OpenStudy (imstuck):

Cool!

OpenStudy (anonymous):

ok i thought i new how but i would need the angles in order to do what i was thinking

OpenStudy (imstuck):

Use Pythagorean's theorem. Yes? You know it?

OpenStudy (anonymous):

oh yes i know that

OpenStudy (anonymous):

2^2-9^2= 85 c^2= 85 c=9.2 ? i think this is right